A solution that is 34% antifreeze is going to be enriched by adding pure antifreeze. How much of each is needed to generate 4 gallons of a solution that is 63% antifreeze?
Found 2 solutions by richwmiller, ankor@dixie-net.com:
Answer by richwmiller(17219)

.34x+y=.63(4),
x+y=4
x=2.24242, y=1.75758

Answer by ankor@dixie-net.com(22740)

A solution that is 34% antifreeze is going to be enriched by adding pure antifreeze. How much of each is needed to generate 4 gallons of a solution that is 63% antifreeze?
:
let x = amt of 34% solution
the total amt is to be 4 gal, therefore
(4-x) = amt of pure antifreeze
:
.34x + (4-x) =.63(4)
.34x + 4 - x = 2.52
-.66x = 2.52 - 4
-.66x = -1.48
x = -1.48/-.66
x = 2.24 gal of the 34% solution
then
4 - 2.24 = 1.76 gal of the pure antifreeze